2 particles carrying charges -q and +q have equal mass m each,is fixed at the ends of a light rod of length a to form a dipole.rod is clamped at an end and is placed in a uniform electric field E with axis of the dipole along the electric field.rod is slightly tilted and then released.neglecting gravity,find time period of oscillations.

Grade:

1 Answers

Ankit Khokhar
12 years ago
Hi swapan!!!!! You have posted a very good question. Lets see....... We know that torque=p*E (i have used * for cross product) i.e. torque=(p)(e)(sin x) (where x is the angle between p and e) This concept is also given in NCERT. Now torque=I(alpha)=(2m)/((12)(l^2))*alpha (here * is only for simple product) now use alpha=wdw/dx and integrate. (w refers to omega angular velocity) after that take w=dx/dt and solve you will get the answer.